Database last updated on Wed, 28 Jan 2026 15:10:56 +0000 by localhost

Instrument constant changes for the week before 2025010 (2 sec, 0 rows)

HG burned out? (AP results when HG is ON) (0 sec, 1 rows)

SerialYYYYJJJHG CurrentHG Voltage
1952025007013.78

Measurements and tests summary.


_jday_loc_serial_skc
010STONY PLAIN#195menu
(0 sec)

Dead time (0 sec)

Nothing to report.

Standard lamp R6 (1 sec)

Nothing to report.

Standard lamp R5 (1 sec)

Nothing to report.

Number of days since last dome/window cleaning (0 sec, 2 rows)

serialdays
195more than a year

Missing files (0 sec, 1 rows)

Seriallast datadays ago
19520250110-1

Ozone results for good DS obs (std<2.5DU, mu<4) within 3 minutes between 2 Brewers. (0 sec)

No ozone measurements to display.

Ozone results for good FM obs (std<5DU, mu<4) within 3 minutes between 2 Brewers. (0 sec)

No ozone measurements to display.

Unsorted comments (1 sec, 9819 rows)

SerialtypeTextCountYYYYJJJ
195env:DosBox12025010
195cubdsp:112025010

The END (0 sec, 2 rows)

If you have questions and/or comments about this page please contact Volodya Savastiouk